# SCMPPI: Supervised Contrastive Multimodal Framework for Predicting Protein-Protein Interactions

### v1

@@ A.2.2 The k-Spaced (k) and similarity threshold ($\tau$) on model
In our experiments using the Human dataset, we optimized two hyperparameters k-Spaced (k) and similarity threshold ($\tau$). Among them, the values of k are in the range of 0, 1, 2, 3, 4, and the experimental results show that the model performs best when k=3 after considering the sensitivity , F1 score, and runtime, so we determined k=3 as the final parameter setting (Equation 5). On the other hand, the $\tau$ was traversed in the range of 0.0, 0.3, 0.5, 0.7, 1.0, and the experimental results showed that the negative sample filter could provide higher sensitivity , lower false-negative rate, and optimal F1 scores when $\tau$=0.7, so we chose $\tau$=0.7 as the final threshold setting (Eq. 8).
@@ B.1 More Baseline Results and Ablation
Table 8: Performance comparison on dataset Human (clc=1.0)
@@
Table 9: Performance comparison on dataset H.pylori (clc=1.0)
@@
Table 10: Performance comparison on dataset pipr-cut (clc=0.6)
@@
Table 11: Performance comparison on multi-species(C. eleg, D. mela and E. coli) dataset (clc=0.3).

### v2

@@ A.2.4 The k-Spaced (k) and similarity threshold ($\tau$) on model
On the other hand, we traversed the range of 0.0, 0.3, 0.5, 0.7, 1.0 to get $\tau$ in Figure 4, and the experimental results showed that the negative sample filter can provide higher sensitivity, lower false negative rate and best F1 score when $\tau$=0.7, so we chose $\tau$=0.7 as the final threshold setting (Equation 7) .
@@ B.1 More Baseline Results and Ablation
Table 5: Performance comparison on dataset Human ($\kappa$=1.0)
@@
Table 6: Performance comparison on dataset H.pylori ($\kappa$=1.0)
@@
Table 7: Performance comparison on dataset pipr-cut ($\kappa$=0.6)
@@
Table 8: Performance comparison on multi-species(C. eleg, D. mela and E. coli) dataset ($\kappa$=0.3).